NI intensity stabilization

Go down

NI intensity stabilization

Post  Alessio on Fri Jan 29, 2010 5:54 pm

(Viviana, Alessio)

This morning we worked on the NI intensity stabilization. We have
set the laser temperature to about 20.9 deg, that gives a
good stability of the output power.
Figure 1 shows a comparison between the open loop noise spectrum
(pink) and closed loop spectrum (black). The reduction factor is the
same as the WI bench (our reference).
We also investigated the origin of the high frequency bumps (from
100Hz on). We compared the closed loop spectrum with the dark noise of
the photodiodes:
1) figure 2: against the out-of-loop PD, it does not seem to be the cause
2) figure 3: against the in-loop PD, it seems to be the cause
3) figure 4: to understand if the culprit is the PD or the low noise
amplifier, we connected the amplifier to the out-of-loop PD. thus it
seems that the source is the pre-amplifier.
We want to underline that, since these bumps are at high frequency, they are not important for the sensitivity.





avatar
Alessio

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum